OpenStudy (anonymous):

Find the vector and parametric quations of the line parallel to [2 -1 0] and passing throught P(1,-1,3)

OpenStudy (anonymous):

quations* equations

OpenStudy (nenadmatematika):

parametric equation is:\[x=2t+1\] ; \[y=-t-1\] ; \[z=3\] and the vector equation is: (x-1)/2=(y+1)/(-1)=(z-3)/0